40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 442     Entry time: Thu Apr 24 14:10:26 2008
Author: rob 
Type: Update 
Category: Locking 
Subject: locking work 
Rob, Johnnie

We made some progress on locking last night (Wed night), namely that we were able to handoff (briefly) the CARM-MCL path the REFL-DC error signal. We tried this because we suspect that the reason the PO-DC is not a good CARM error signal is because at low powers, the dc light level in the recycling cavity is dominated by the +f2 RF sideband. Thus, REFL-DC should work a bit better at low powers, which it did. It wasn't super stable, though, so this will require a bit of work to make the transition reliable & stable. The next things to work on include setting the AO path gain properly and possibly going to higher arm powers before handing off (thus increasing the discriminant).

Another thing we found is that the alignment scripts are not working in an ideal fashion. Running the alignment scripts for the two arms (XARM & YARM) leaves the Michelson badly misaligned, making it impossible to get good DRM alignment. This will have to be fixed.
ELOG V3.1.3-